A beaker contains a solution that is 15% alcohol. In another beaker is a solution that is 40% alcohol. How mcuh of each should be mixed to obtain 9 L of solution that is 25% alcohol?
Answer by checkley77(12844)

.40x+.15(9-x)=.25*9
.40x+1.35-.15x=2.25
.25x=2.25-1.35
.25x=.9
x=.9/.25
x=3.6 L of 40% solution is used.
9-3.6=5.4 L of 15% solution is used.
Proof:
.40*3.6+.15*5.4=2.25
1.44+.81=2.25
2.25=2.25